### v1

@@ Abstract
[1]organization=Indian Institute of Space Science and Technology, addressline=Valiamala Road, Valiamala, city=Thiruvananthapuram, postcode=695547, country=India
@@
[2]organization=Atria University, addressline=1st Main Road, AGS Colony, Anandnagar, Hebbal, city=Bengaluru, postcode=560024, country=India
@@
[3]organization=Indian Institute of Science, addressline=CV Raman Road, city=Bengaluru, postcode=560012, country=India
@@
[4]organization=Vikram Sarabhai Space Centre, addressline=Perumathura Road, Veli, city=Thiruvananthapuram, postcode=695022, country=India
@@
[5]organization=Space Applications Centre, addressline=Satellite Road, city=Ahmedabad, postcode=380015, country=India
@@ 1 Introduction
The quest to understand the origins of life on Earth and its possibility elsewhere in the universe has become a frontier theme in humanity’s exploration of outer space. Major space organizations, including the Indian Space Research Organisation (ISRO) with its planned Gaganyaan human space flight program [^2] https://www.isro.gov.in/Gaganyaan.html, continue to explore in various experimental ways the impact of the hostile conditions that prevail in space (microgravity, ultraviolet flux, and particle radiation field) on biological systems. Understanding the impact of space on microbial life forms has been a dominant theme of research in this direction with several objectives and approaches as summarised in Horneck G [11]. In this context, the Small-spacecraft Systems and PAyload CEntre [^3] https://www.iist.ac.in/sspace (SSPACE) at the Indian Institute of Space Science and Technology (IIST), Thiruvananthapuram has initiated the SSPACE Astrobiology Payload program (SAP), with SAP-1 as its first mission. This initiative complements missions like ISRO’s Gaganyaan, India’s ambitious manned space mission, by providing essential insights into how space travel affects biological entities, a critical aspect for the success and safety of manned missions.
@@
The SAP is a long-term program aimed at developing modular devices for conducting in-situ microbiology experiments in the low earth orbit (LEO). While drawing parallels with NASA’s Organism/Organic Exposure to Orbital Stresses (OOREOS) satellite [8], SAP’s primary objective is to create versatile modules adaptable for small satellites or space platforms aligned with ISRO’s missions. The genesis of this program was inspired by the proposed microgravity experiment [12], Microbial Analysis in Space (MANAS) which is focused on studying the growth patterns of Sporosarcina pasteurii, a spore-forming bacterium renowned for its ability to induce calcite precipitation. This foundation paves the way for SAP-1, the initial mission in a series of astrobiology research endeavors by the SSPACE at IIST. SAP-1 and its successors aim to uncover astrobiological insights and significantly contribute to the success of future human space exploration, leveraging small satellites for a broad spectrum of microbiological studies in LEO.
@@
SAP-1 aims to study the intricate interactions between microorganisms and the harsh conditions of space, particularly microgravity, by employing optical density measurement techniques. It focuses on the growth characteristics of probiotic bacteria, specifically Bacillus clausii and Bacillus coagulans, which have shown significant health benefits but remain relatively unexplored in the context of adverse space conditions of the Low Earth Orbit (LEO). The objectives of SAP-1 extend beyond theoretical exploration, aiming to provide tangible insights that can enhance the safety and success of future human space missions. The SAP-1 platform addresses the multifaceted challenges posed to microbiology experiments by outer space conditions, including maintaining optimal pressure and temperature, and supplying essential nutrients for bacterial growth. To realize these goals, a novel approach employing a rotating platform with transparent chambers (bio-wells) and a nutrition media reserve has been devised, offering a comprehensive solution to the critical challenges.
@@ 2.1 Introduction
Microbiological experiments conducted in LEO serve as pivotal endeavours in advancing our comprehension of how living organisms adapt to the formidable challenges presented by the space environment. LEO is distinguished by a confluence of particle radiations emanating from galactic and solar sources, solar electromagnetic radiations at ultraviolet energies, the absence of atmospheric pressure, substantial temperature oscillations, and microgravity, collectively rendering it an inhospitable milieu for biological systems. The exposure to microgravity, along with ionizing radiation originating from both solar and galactic origins, leads to an adverse space environment detrimental to biological entities. The objectives of space biology experiments encompass the provision of empirical evidence regarding the enduring impacts of space on life systems, as well as the exploration of the broader inquiry into the origins and proliferation of life beyond Earth’s confines. The former objective holds direct relevance for both short and long-term human spaceflight programs, as the challenging extraterrestrial conditions expose astronauts to multifaceted physiological and psychological threats [2; 3; 7]. The influences of reduced gravity and ionizing radiations extend to sub-cellular levels, perturbing gene expression and cellular growth pathways [2].
@@ 2.2 Gut Microbiota
The human digestive tract harbors a diverse community of over a trillion beneficial microorganisms, each individual exhibiting a unique variance and abundance of these entities, thereby playing integral roles in metabolism, immunity, and nutrition [26; 5]. Research indicates that spaceflight induces compositional alterations in the microbiota of the gastrointestinal tract, skin, nasal passages, and oral cavity, consequently impacting diverse biological functions [28]. The gut microbiome significantly contributes to vital physiological processes, including digestion, immune system development, central nervous system maturation, and vitamin production [6; 27].
Moreover, the health of the gut intricately influences the emotional and cognitive centers of the brain through the communication pathways of the nervous system, collectively known as the gut-brain axis. Extended periods in space can impose stress on gut homeostasis, the capacity to adapt to changing external conditions, thereby affecting both physiological and psychological well-being. To address such challenges, the oral administration of probiotics emerges as a validated approach for restoring balance to the gut microbiome, enhancing immune function, and sustaining gut health [6].
The potential advantages presented by probiotics underscore the need for extensive research in this relatively unexplored domain. Probiotics, characterized by their diverse benefits and facile administration, emerge as ideal health supplements for astronauts. However, the viability and efficacy of probiotic samples when stored in the space environment for extended duration remain a relatively unexplored aspect, warranting further investigation.
@@ 2.3 Probiotics in space
Current research provides insights into the pivotal role of gut microbiota in modulating the gut-brain axis, a bidirectional communication network that establishes a connection between the central nervous system and the enteric nervous system (ENS). This intricate nexus integrates cognitive and emotional brain centers with peripheral intestinal operations, allowing for bidirectional interplay between the microbiota and the gut-brain axis through signaling mechanisms extending from the gut microbiota to the central nervous system and vice versa [4].
@@
The homeostasis of the gut microbiota, crucial for efficient communication between the brain and gut, can be influenced by the space environment, potentially disrupting optimal gut-brain communication. To address this, the administration of probiotics, known for their capacity to harmonize the gut microbiome, emerges as an effective strategy [22].
@@
Various factors in the space environment, including microgravity, seclusion, circadian rhythm changes, and additional stressors, can precipitate psychological disturbances in astronauts, such as anxiety, depression, and related mental health challenges [19; 23]. Emotional well-being is as paramount as physical well-being for astronauts conducting experiments in space. Disruption in microbiota homeostasis can impact mental health by directly influencing the release of neurotransmitters such as serotonin and dopamine, regulating the stress response by affecting the hypothalamus–pituitary–adrenal axis. Microbiota alterations can also influence the levels of BDNF (brain-derived neurotrophic factor) and stimulate inflammatory cytokine release, contributing to the intricate interplay between gut health and mental well-being [1]. Empirical research indicates an increased propensity for emotional responses and psychiatric disorders in astronauts during space missions. Clinical investigations, including controlled trials and observational studies on probiotics’ application for anxiety and depression management, have demonstrated a reduction in the severity of these neuropsychiatric conditions [1; 16]. Hence, probiotics have the potential to contribute significantly to the mental well-being of space travelers.
@@
In the space environment, astronauts encounter gastrointestinal disturbances, including epithelial barrier disruption and immunological dysregulation, potentially attributed to microgravity’s effects on the intestinal epithelium [14]. Studies indicate [2] an increase in pathogenic potential and antibiotic resistance in specific microorganisms following spaceflight exposure. Certain strains of probiotic bacteria exhibit antimicrobial characteristics, often linked to the secretion of specific peptides or bioactive compounds. These molecules not only confer a competitive advantage within the intricate gut microbiome but also provide host protection against pathogenic bacteria, thereby promoting the persistence of commensal bacteria. Additionally, certain probiotic bacteria strains have been demonstrated to modulate mucin expression, consequently altering the mucus barrier’s characteristics and indirectly influencing the gastrointestinal immune response [2].
@@ 2.4 Bacillus clausii
Upper respiratory tract infections stand among the most prevalent infections in pediatric populations. Research suggests positive outcomes associated with probiotics in treating these infections, particularly the efficacy of Bacillus clausii UBBC-07 in managing upper respiratory tract infection symptoms [18; 9]. A key rationale for administering Bacillus clausii as a probiotic during antibiotic regimens lies in its intrinsic resilience to a broad spectrum of antibiotics. Consequently, it retains the ability to populate the gastrointestinal tract even in the presence of reduced commensal bacteria resulting from antibiotic exposure, thus aiding in mitigating the risk of gut dysbiosis. Bacillus clausii secretes compounds with antimicrobial activity against diverse pathogens, playing a pivotal role in maintaining a harmonious gastrointestinal microbial equilibrium [17; 29].
The emerging consensus on the significant role of gut microbiota in modulating responses to cancer treatments has prompted attention. A robust and diverse gut microbial composition is believed to enhance outcomes of immunotherapy and chemotherapy [10; 20]. Many gut bacteria, including Bacillus clausii, produce metabolites with observed anti-tumorigenic properties. Bacillus clausii is part of a distinct intra-tumoral microbial signature associated with prolonged survival in patients with surgically resected pancreatic adenocarcinoma, along with Pseudoxanthomonas, Streptomyces, and Saccharopolyspora [25]. The spores of Bacillus clausii UBBC-07 have also demonstrated benefits in ameliorating oral mucositis symptoms in patients with head and neck cancer undergoing radiation treatment [21].
Recognition of the gut-brain axis and the potential therapeutic role of probiotics is growing for enhancing interventions in stress and various neurological conditions. Bacillus clausii has demonstrated the ability to alleviate stress-related behavior in mice through augmentation of monoamine levels, along with an increase in the mRNA expression of dopamine receptors (D1 and D2) and synaptophysin [24]. Despite the essential role of riboflavin in cellular functioning and growth, humans do not endogenously synthesize it. Specific strains of Bacillus clausii can bio-synthesize and secrete riboflavin, potentially contributing to enhanced cellular functionality and growth in humans [9]. A comprehensive analysis of the entire genome sequence of Bacillus clausii UBBC07 has revealed the presence of antibiotic-resistance genes within its chromosomal DNA, which are intrinsic and not transferable. Additionally, no toxin genes were detected, collectively indicating that the consumption of Bacillus clausii UBBC07 is considered safe for human consumption [15].
@@ 2.5 SAP-1 payload with Bacillus clausii
Fig. 2: A flowchart of order of events to be conducted onboard SAP-1 during the microbiology experiment (image generated using bioRender).
Utilizing the SAP-1 bio hardware developed by SSPACE at IIST, dry bacterial spores can be transported to space in small bio-wells, accompanied by the provision of supplying nutrient broth at the designated activation time (Figure 2). Growth rate tracking, facilitated by optical density measurements, is a valuable tool for comparing space-based outcomes with ground-based control experiments and deciphering microbial growth in the space environment. The storage of Bacillus clausii spores at 30°C, followed by incubation at 35°C for germination and culturing, is the proposed methodology. Once the culture has saturated and reached the senescence phase, the experiment will be terminated by heating the payload to over 120°C, thus preventing any kind of biological contamination in space. Once the culture has saturated and reached the senescence phase, the experiment will be concluded by elevating the payload’s temperature to over 120 degrees Celsius. Similar measures are employed in the event of an intermediate termination of the experiment. This is implemented to prevent any potential biological contamination in the space environment effectively.
@@ 2.6 Experiments to determine optimal aeration
Bio-wells on SAP-1 are air-tight chambers and hence the apt ratio of the quantity of atmospheric air to the quantity of bacterial culture must be known for optimal growth of bacteria. An experiment was carried out using Bacillus clausii and Bacillus coagulans to determine a favourable ratio by varying aeration levels and quantity of strain cultures and utilizing a closed 24-well plate and falcon tubes for inoculating different culture ratios, monitoring and recording optical density measurements over a defined period.
In the experimental methodology, cultures were cultivated in standard Nutrient Broth. The preparation of cultures involved utilizing a 5 mg/ml spore suspension as an inoculum, with inoculation performed using a 10% suspension of the final volume. The varied volumes of culture across samples were set at 25% (600 ul), 50% (1200 ul), and 95% (2400 ul) of the total well volume, representing 75%, 50%, and 5% of air volume, respectively. Triplicate samples were grown. Cultures were incubated in a multi-well plate reader for 16 hours at 37℃ and 180 rpm. Optical density measurements at 600 nm were taken every 30 minutes.
To address path length discrepancies resulting from different culture volumes, a calibration step was introduced. This involved measuring the optical density of equivalent volumes from a single saturated culture. Since optical density was not directly proportional to sample volume (path length), actual OD readings were normalized using the corresponding calibration factor.
@@
Our experiments indicate that an aeration volume of 75% is optimal for obtaining Optical Density (OD) measurements over a relatively extended period.

### v2

@@ Abstract
[1]organization=Indian Institute of Space Science and Technology, addressline=Valiamala Road, Valiamala, city=Thiruvananthapuram, Kerala, postcode=695547, country=India
@@
[2]organization=Department of Physics, Indian Institute of Technology Tirupati, addressline=Venkatagiri road, Yerpedu, city=Tirupati, Andhra Pradesh, postcode=517619, country=India
@@
[3]organization=Atria University, addressline=1st Main Road, AGS Colony, Anandnagar, Hebbal, city=Bengaluru, Karnataka, postcode=560024, country=India
@@
[4]organization=Indian Institute of Science, addressline=CV Raman Road, city=Bengaluru, Karnataka, postcode=560012, country=India
@@
[5]organization=Vikram Sarabhai Space Centre, addressline=Perumathura Road, Veli, city=Thiruvananthapuram, Kerala, postcode=695022, country=India
[…]
[6]organization=Space Applications Centre, addressline=Satellite Road, city=Ahmedabad, Gujarat, postcode=380015, country=India
@@ 1 Introduction
The quest to understand the origins of life on Earth and its possibility elsewhere in the universe has become a frontier theme in humanity’s exploration of outer space. Major space organizations, including the Indian Space Research Organisation (ISRO) with its planned Gaganyaan human space flight program [^2] https://www.isro.gov.in/Gaganyaan.html, continue to explore in various experimental ways the impact of the hostile conditions that prevail in space (microgravity, ultraviolet flux, and particle radiation field) on biological systems. Understanding the impact of space on microbial life forms has been a dominant theme of research in this direction with several objectives and approaches as summarised in Gerda et al. (2010). In this context, the Small-spacecraft Systems and PAyload CEntre [^3] https://www.iist.ac.in/sspace (SSPACE) at the Indian Institute of Space Science and Technology (IIST), Thiruvananthapuram has initiated the SSPACE Astrobiology Payload program (SAP), with SAP-1 as its first mission. This initiative complements missions like ISRO’s Gaganyaan, India’s ambitious manned space mission, by providing essential insights into how space travel affects biological entities, a critical aspect for the success and safety of manned missions.
@@
The SAP is a long-term program aimed at developing modular devices for conducting in-situ microbiology experiments in the low earth orbit (LEO). While drawing parallels with NASA’s Organism/Organic Exposure to Orbital Stresses (OOREOS) satellite (Ehrenfreund et al., 2014), SAP’s primary objective is to create versatile modules adaptable for small satellites or space platforms aligned with ISRO’s missions. The genesis of this program was inspired by the proposed microgravity experiment (Kallapur et al., 2021a), Microbial Analysis in Space (MANAS) which is focused on studying the growth patterns of Sporosarcina pasteurii, a spore-forming bacterium renowned for its ability to induce calcite precipitation. This foundation paves the way for SAP-1, the initial mission in a series of astrobiology research endeavors by the SSPACE at IIST. SAP-1 and its successors aim to uncover astrobiological insights and significantly contribute to the success of future human space exploration, leveraging small satellites for a broad spectrum of microbiological studies in LEO.
@@
SAP-1 aims to study the intricate interactions between microorganisms and the harsh conditions of space, particularly microgravity, by employing optical density (OD) measurement techniques. It focuses on the growth characteristics of probiotic bacteria, specifically Bacillus clausii and Bacillus coagulans, which have shown significant health benefits but remain relatively unexplored in the context of adverse space conditions of the LEO. The objectives of SAP-1 extend beyond theoretical exploration, aiming to provide tangible insights that can enhance the safety and success of future human space missions. The SAP-1 platform addresses the multifaceted challenges posed to microbiology experiments by outer space conditions, including maintaining optimal pressure and temperature and supplying essential nutrients for bacterial growth. To realize these goals, a novel approach employing a rotating platform with transparent chambers (bio-wells) and a nutrition media reserve has been devised, offering a comprehensive solution to the critical challenges.
@@ 2.1 Introduction
Microbiological experiments conducted in LEO serve as pivotal endeavours in advancing our comprehension of how living organisms adapt to the formidable challenges presented by the space environment. LEO is distinguished by a confluence of particle radiations emanating from galactic and solar sources, solar electromagnetic radiations at ultraviolet energies, the absence of atmospheric pressure, substantial temperature oscillations, and microgravity, collectively rendering it an inhospitable milieu for biological systems. The exposure to microgravity, along with ionizing radiation originating from both solar and galactic origins, leads to an adverse space environment detrimental to biological entities. The objectives of space biology experiments encompass the provision of empirical evidence regarding the enduring impacts of space on life systems, as well as the exploration of the broader inquiry into the origins and proliferation of life beyond Earth’s confines. The former objective holds direct relevance for both short and long-term human spaceflight programs, as the challenging extraterrestrial conditions expose astronauts to multifaceted physiological and psychological threats (Bharindwal et al., 2023; Blaber et al., 2010; Crucian et al., 2016). The influences of reduced gravity and ionizing radiations extend to sub-cellular levels, perturbing gene expression and cellular growth pathways (Bharindwal et al., 2023).
@@ 2.2 Gut Microbiota
The human digestive tract harbors a diverse community of over a trillion beneficial microorganisms, each individual exhibiting a unique variance and abundance of these entities, thereby playing integral roles in metabolism, immunity, and nutrition (Sidhu & van der Poorten, 2017; Consortium, 2012). Research indicates that spaceflight induces compositional alterations in the microbiota of the gastrointestinal tract, skin, nasal passages, and oral cavity, consequently impacting diverse biological functions (Voorhies et al., 2019). The gut microbiome significantly contributes to vital physiological processes, including digestion, immune system development, central nervous system maturation, and vitamin production (Cresci & Bawden, 2015; Sorboni et al., 2022).
Moreover, the health of the gut intricately influences the emotional and cognitive centers of the brain through the communication pathways of the nervous system, collectively known as the gut-brain axis (see e.g. Carabotti et al., 2015b). Extended periods in space can impose stress on gut homeostasis, the capacity to adapt to changing external conditions, thereby affecting both physiological and psychological well-being. To address such challenges, the oral administration of probiotics emerges as a validated approach for restoring balance to the gut microbiome, enhancing immune function, and sustaining gut health (Cresci & Bawden, 2015).
The potential advantages presented by probiotics underscore the need for extensive research in this relatively unexplored domain. Probiotics, characterized by their diverse benefits and facile administration, emerge as ideal health supplements for astronauts. However, the viability and efficacy of probiotic samples when stored in the space environment for an extended duration remain a relatively unexplored aspect, warranting further investigation.
@@ 2.3 Probiotics in space
Current research provides insights into the pivotal role of gut microbiota in modulating the gut-brain axis, a bidirectional communication network that establishes a connection between the central nervous system and the enteric nervous system (ENS). This intricate nexus integrates cognitive and emotional brain centers with peripheral intestinal operations, allowing for bidirectional interplay between the microbiota and the gut-brain axis through signaling mechanisms extending from the gut microbiota to the central nervous system and vice versa (Carabotti et al., 2015a).
@@
The homeostasis of the gut microbiota, crucial for efficient communication between the brain and gut, can be influenced by the space environment, potentially disrupting optimal gut-brain communication. To address this, the administration of probiotics, known for their capacity to harmonize the gut microbiome, emerges as an effective strategy Nobile et al. (2021).
@@
Various factors in the space environment, including microgravity, seclusion, circadian rhythm changes, and additional stressors, can precipitate psychological disturbances in astronauts, such as anxiety, depression, and related mental health challenges (Mallis & DeRoshia, 2005; Oluwafemi et al., 2021). Emotional well-being is as paramount as physical well-being for astronauts conducting experiments in space. Disruption in microbiota homeostasis can impact mental health by directly influencing the release of neurotransmitters such as serotonin and dopamine, regulating the stress response by affecting the hypothalamus–pituitary–adrenal axis. Microbiota alterations can also influence the levels of BDNF (brain-derived neurotrophic factor) and stimulate inflammatory cytokine release, contributing to the intricate interplay between gut health and mental well-being (Alli et al., 2022). Empirical research indicates an increased propensity for emotional responses and psychiatric disorders in astronauts during space missions. Clinical investigations, including controlled trials and observational studies on probiotics’ application for anxiety and depression management, have demonstrated a reduction in the severity of these neuropsychiatric conditions (Alli et al., 2022; Liu et al., 2019). Hence, probiotics have the potential to contribute significantly to the mental well-being of space travelers.
@@
In the space environment, astronauts encounter gastrointestinal disturbances, including epithelial barrier disruption and immunological dysregulation, potentially attributed to microgravity’s effects on the intestinal epithelium (La Fata et al., 2018). Studies indicate (Bharindwal et al., 2023) an increase in pathogenic potential and antibiotic resistance in specific microorganisms following spaceflight exposure. Certain strains of probiotic bacteria exhibit antimicrobial characteristics, often linked to the secretion of specific peptides or bioactive compounds. These molecules not only confer a competitive advantage within the intricate gut microbiome but also provide host protection against pathogenic bacteria, thereby promoting the persistence of commensal bacteria. Additionally, certain probiotic bacteria strains have been demonstrated to modulate mucin expression, consequently altering the mucus barrier’s characteristics and indirectly influencing the gastrointestinal immune response (Bharindwal et al., 2023).
@@ 2.4 Bacillus clausii
Upper respiratory tract infections stand among the most prevalent infections in pediatric populations. Research suggests positive outcomes associated with probiotics in treating these infections, particularly the efficacy of Bacillus clausii UBBC-07 in managing upper respiratory tract infection symptoms Madempudi et al. (2022); Ghelardi et al. (2022). A key rationale for administering Bacillus clausii as a probiotic during antibiotic regimens lies in its intrinsic resilience to a broad spectrum of antibiotics. Consequently, it retains the ability to populate the gastrointestinal tract even in the presence of reduced commensal bacteria resulting from antibiotic exposure, thus aiding in mitigating the risk of gut dysbiosis. Bacillus clausii secretes compounds with antimicrobial activity against diverse pathogens, playing a pivotal role in maintaining a harmonious gastrointestinal microbial equilibrium Lopetuso et al. (2016); Wong-Chew et al. (2022).
The emerging consensus on the significant role of gut microbiota in modulating responses to cancer treatments has prompted attention. A robust and diverse gut microbial composition is believed to enhance outcomes of immunotherapy and chemotherapy Gopalakrishnan et al. (2018); Matson et al. (2018). Many gut bacteria, including Bacillus clausii, produce metabolites with observed anti-tumorigenic properties. Bacillus clausii is part of a distinct intra-tumoral microbial signature associated with prolonged survival in patients with surgically resected pancreatic adenocarcinoma, along with Pseudoxanthomonas, Streptomyces, and Saccharopolyspora Riquelme et al. (2019). The spores of Bacillus clausii UBBC-07 have also demonstrated benefits in ameliorating oral mucositis symptoms in patients with head and neck cancer undergoing radiation treatment Mirza et al. (2022).
Recognition of the gut-brain axis and the potential therapeutic role of probiotics is growing for enhancing interventions in stress and various neurological conditions. Bacillus clausii has demonstrated the ability to alleviate stress-related behavior in mice through augmentation of monoamine levels, along with an increase in the mRNA expression of dopamine receptors (D1 and D2) and synaptophysin Rehman et al. (2022). Despite the essential role of riboflavin in cellular functioning and growth, humans do not endogenously synthesize it. Specific strains of Bacillus clausii can bio-synthesize and secrete riboflavin, potentially contributing to enhanced cellular functionality and growth in humans Ghelardi et al. (2022). A comprehensive analysis of the entire genome sequence of Bacillus clausii UBBC07 has revealed the presence of antibiotic-resistance genes within its chromosomal DNA, which are intrinsic and not transferable. Additionally, no toxin genes were detected, collectively indicating that the consumption of Bacillus clausii UBBC07 is considered safe for human consumption Lakshmi et al. (2017).
@@ 2.5 SAP-1 payload with Bacillus clausii
Fig. 2: A flowchart of the order of events to be conducted onboard SAP-1 during the microbiology experiment (image generated using bioRender).
Utilizing the SAP-1 bio hardware developed by SSPACE at IIST, dry bacterial spores can be transported to space in small bio-wells, accompanied by the provision of supplying nutrient broth at the designated activation time (Figure 2). The current version of the SAP-1 has four bio-wells, however, this number can be easily increased in the final space-grade version. Growth rate tracking, facilitated by OD measurements, is a valuable tool for comparing space-based outcomes with ground-based control experiments and deciphering microbial growth in the space environment. The storage of Bacillus clausii spores at 30°C, followed by incubation at 35°C for germination and culturing, is the proposed methodology. Once the culture has saturated and reached the senescence phase, the experiment will be terminated by heating the payload to over 120°C, thus preventing any kind of biological contamination in space. Once the culture has saturated and reached the senescence phase, the experiment will be concluded by elevating the payload’s temperature to over 120 degrees Celsius. Similar measures are employed in the event of an intermediate termination of the experiment. This is implemented to prevent any potential biological contamination in the space environment effectively.
@@ 2.6 Experiments to determine optimal aeration
Bio-wells on SAP-1 are air-tight chambers and hence the apt ratio of the quantity of atmospheric air to the quantity of bacterial culture must be known for optimal growth of bacteria. An experiment was carried out using Bacillus clausii and Bacillus coagulans to determine a favourable ratio by varying aeration levels and quantity of strain cultures and utilizing a closed 24-well plate and falcon tubes for inoculating different culture ratios, monitoring and recording OD measurements over a defined period.
In the experimental methodology, cultures were cultivated in standard Nutrient Broth. The preparation of cultures involved utilizing a 5 mg/ml spore suspension as an inoculum, with inoculation performed using a 10% suspension of the final volume. The varied volumes of culture across samples were set at 25% (600 ul), 50% (1200 ul), and 95% (2300 ul) of the total well volume, representing 75%, 50%, and 5% of air volume, respectively. Triplicate samples were grown. Cultures were incubated in a multi-well plate reader for 16 hours at 37℃ and 180 rpm. OD measurements at 600 nm were taken every 30 minutes.
To address path length discrepancies resulting from different culture volumes, a calibration step was introduced. This involved measuring the OD of equivalent volumes from a single saturated culture. Since OD was not directly proportional to sample volume (path length), actual OD readings were normalized using the corresponding calibration factor.
@@
Our experiments indicate that an aeration volume of 75% is optimal for obtaining OD measurements over a relatively extended period. While this 75% optimal aeration volume is based on ground tests, microgravity may slightly alter the required volume due to changes in fluid dynamics - specifically the absence of gravitational convection. However, the chosen volume is sufficient to meet the needs of the initial bacteria and nutrients.